OpenStudy (daniel.ohearn1):

Integrate (x/(x-6))dx...

OpenStudy (irishboy123):

hint \(\dfrac{x}{x + a} = \dfrac{x+ a - a}{x + a} \)

OpenStudy (daniel.ohearn1):

How about the Integral of (x-9)/(x+5)(x-2) ?

OpenStudy (irishboy123):

post it as a new quessie? you can still kick off with the same idea such as \(\dfrac{x-9}{(x+5)(x-2)} = \dfrac{x-2 - 7}{(x+5)(x-2)} \) you will still need to know how to decompose, though